OpenStudy (anonymous):

an architect designed an outdoor staircase for a house.the relationship between the height of the steps and the length of the tread is modeled b y the equation 57x-95y=0 which of the following represents the slope 5/3 3/2 2/3 3/5

OpenStudy (amistre64):

oy.... i built hundreds of staircases as a carpenter

OpenStudy (amistre64):

slope = -a/b -57/-95

OpenStudy (amistre64):

im guessing 3/5

OpenStudy (amistre64):

19*3 = 27+30 = 57 19*5 = 50 +45 = 95

OpenStudy (anonymous):

thank you